Find the average value (g ave) of the function g on the given interval.
g(x) = 2*​x^​(1/​3) [8,125]
Answer #1

Sol glas Given function gla) = 243 The average value of function on the interval [a,b ] is gave a g (4 dn 125 La 2x 3 che 8 3
Answer #2

SOLUTION :


Average value (g.)  of the given function g(x) in interval (1, 27)


   125

= ∫    g(x) dx  / (125 - 8)

   8


              125

= 1/117   ∫    2 (x^(1/3)  dx

               8

                                 125

= 2/117 * 3/4 [ x^4/3) ]                            

                                  8                                                      


= 3/234 * [125^(4/3) - 8^(4/3)]


= 3/234 * 609


= 7.807  (ANSWER)

answered by: Tulsiram Garg